Solve by cauchy integral formula
∫Z^2 e^z dz ; C : |z| = 1
1
Expert's answer
2020-10-26T19:54:41-0400

Сauchy integral formula: "f(z_0) = \frac{1}{2\pi i}\int_c \frac{f(z)}{(z-z_0)}"

"\int_c z^2 e^zdz = \int_c \frac{z^3e^z}{z}dz \\ f(z) = z^3 e^z, z_0 = 0"

So "\int_cz^2e^z = 2\pi i f(z_0) = 2\pi i *0*e^0 = 0"